# Online Optimization for Learning to Communicate over Time-Correlated Channels

### v4

@@ Abstract
Machine learning techniques have garnered great interest in designing communication systems owing to their capacity in tacking with channel uncertainty. To provide theoretical guarantees for learning-based communication systems, some recent works analyze generalization bounds for devised methods based on the assumption of Independently and Identically Distributed (I.I.D.) channels, a condition rarely met in practical scenarios. In this paper, we drop the I.I.D. channel assumption and study an online optimization problem of learning to communicate over time-correlated channels. To address this issue, we further focus on two specific tasks: optimizing channel decoders for time-correlated fading channels and selecting optimal codebooks for time-correlated additive noise channels. For utilizing temporal dependence of considered channels to better learn communication systems, we develop two online optimization algorithms based on the optimistic online mirror descent framework. Furthermore, we provide theoretical guarantees for proposed algorithms via deriving sub-linear regret bound on the expected error probability of learned systems. Extensive simulation experiments have been conducted to validate that our presented approaches can leverage the channel correlation to achieve a lower average symbol error rate compared to baseline methods, consistent with our theoretical findings.
@@ I Introduction
Most existing studies utilizing machine learning approaches to design communication systems lack theoretical justification for their proposed methods, and typically regard learned communication systems as black boxes [5, 6, 7]. Recently, a few works are aiming to conduct theoretical analyses for learning-based communication systems [8, 9, 10]. Specifically, these studies leverage statistical learning theory [11, 12] to derive generalization bounds based on the Independently and Identically Distributed (I.I.D.) channel assumption. Building upon this assumption, the generalization capacity of learning-based communication systems can be assured. This is because these systems can be trained using sufficient data during the offline stage and demonstrate commendable performance when evaluated with data sampled from the same distribution.
However, meeting the assumption of I.I.D. channels in reality is generally challenging, given that practical communication scenarios often involve time-varying channels with dynamic statistical properties. Among time-varying channels, a prominent example is the time-correlated channel [13]. In real-world communication processes, there exist numerous instances of time-correlated channels. For example, user mobility usually leads to time-correlated fading channels in mobile communication [14, 15]. As a consequence of moderate user mobility, channel fading gains become interdependent between consecutive time slots, resulting in the well-known Markov fading channels [16, 17]. Hence, establishing generalization bounds for learning-based communication systems over time-correlated channels is challenging, as statistical learning theory, which depends on the I.I.D. channel assumption, is not applicable in this scenario.
@@
In this paper, we will establish online optimization algorithms [18, 19] for learning communication systems over time-correlated channels with solid theoretical guarantees. Specifically, we consider optimizing the uplink communication process between a mobile user equipment acting as the transmitter and a base station serving as a receiver. One potential application of proposed algorithms is to enhance the performance of distributed learning conducted by multiple mobile devices within wireless networks [20, 21]. In each communication round, we consider that the transmitter sends a package to the receiver. This package is composed of two parts: the first part is a deterministic training sequence known to both the transmitter and receiver, and the second part consists of data symbols that carry the actual information intended for transmission. These data symbols are random and unknown to the receiver. We highlight that this training sequence serves a similar role of pilot signals used in channel estimation in mobile communication, which is essential for calculating the loss required for conducting online optimization. The outlined online optimization procedure is depicted in Fig. 1. Following the assumption widely adopted in the study on using online learning to optimize communication systems [22, 23], we presume that the channel remains constant during the transmission of both the training sequence and the data symbols within a single communication round. While the channel statistics varies across different rounds. Consequently, channel decoders and codebooks learned by training sequences can be utilized to improve the transmission performance of data symbols within the same round. The transmission process of uplink communication is shown in Fig. 2.
@@
For the time-correlated channels studied in this paper, to be specific, we focus on two types of channels: 1) time-correlated fading channel with additive white Gaussian noise, where the fading distribution is unknown and time-correlated; and 2) time-correlated additive noise channel with an unknown and time-correlated noise distribution. For the time-correlated fading channel, the transceiver is equipped with a fixed codebook (constellation) and is tasked with optimizing the channel decoder using collected channel input-output pairs data. Besides, based on the proposed convex surrogate loss, we can regard such online channel decoder learning problem as an online convex optimization problem. In the case of time-correlated additive noise channels, the goal of transceiver becomes to select optimal codebooks during the whole communication process. Similar to [8], the codebook considered in this paper essentially refers to a modulation scheme, i.e., a constellation. We model this online codebook learning problem as a multi-armed bandit problem: the transceiver selects optimal codebooks from a predefined super-codebook using empirical data collected in real-time.
@@ II Related works
In this section, we begin by presenting recent studies that have utilized machine learning techniques in the design of communication systems. Next, we provide an overview of research in online optimization, with an emphasiss on the application of optimistic OMD on predictable benign environments.
As previously mentioned, the remarkable success of machine learning algorithms has spurred interest in their application to optimize communication systems [5]. Some endeavors intend to replace existing components of communication systems with modules learned by empirical data [27]. Another method involves adapting conventional algorithms by integrating deep neural networks [28]. On the other hand, some recent works leverage statistical learning theory for establishing generalization bounds of learned communication systems. For instance, [8] utilize channel input-output pair data to optimize channel decoders and constellations, and derive generalization bounds for learned communication schemes using Rademacher complexity. Besides, [9, 10] employ the probably approximately correct (PAC) learning framework to provide theoretical guarantees for learned communication systems under discrete memoryless channels. Nonetheless, previous works have predominantly conducted a theoretical analysis based on the I.I.D. channel assumption, which is not applicable to with most practical communication scenarios.
Online optimization can effectively model numerous online machine learning problems, encompassing online convex optimization [18], prediction with expert advice [29], multi-armed bandit [30, 31] and more. This framework deals with a sequential decision problem, where a learner repeatedly takes actions within a feasible set and encounters a potentially adversarial loss function from the environment [18, 19]. During this decision-making process, the learner endeavors to devise an algorithm that minimizes regret, defined as the disparity between the total loss incurred by the learner and that of the best decision in hindsight. Departing from statistical learning theory, online optimization theory can furnish theoretical guarantees for algorithms without relying on the I.I.D. assumption. Recently, it has been noted that practical environments often exhibit predictable patterns, which the learner can leverage to achieve reduced regret [24, 26]. To pursue this objective, an optimistic Online Mirror Descent (OMD) framework has been proposed to harness environmental correlations for improving the performance of learned decisions [24, 26].
@@ III-B Time-varying Channel Models
In this paper, we focus on two specific cases of this time-varying channel model, both of which hold high relevance in realistic scenarios. The first case involves a time-correlated fading channel, assuming that $f_{t}$ is a time-correlated linear transformation while the channel noise $Z_{t}$ is an IID Gaussian noise. The second case pertains to a time-correlated additive noise channel, where $f_{t}$ is an identity transformation and $Z_{t}$ is a time-correlated additive channel noise.
@@ IV-A Preliminaries
where $r\geq 1$. (a) holds based on the Boole’s inequality, and (b) follows from the fact that $\mathbb{E}[\mathbbm{1}\{\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}-\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\leq 0\}]\leq\mathbb{E}\big{[}r-\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\big{]}_{+}$ when $r$ satisfies $r\geq 1$.
@@ IV-B Optimistic OMD with Euclidean Regularizer
where $\mathcal{B}_{\psi}(X,Y)=\psi(X)-\psi(Y)-\langle\nabla\psi(Y),X-Y\rangle$ denotes the Bregman divergence induced by a differentiable convex function $\psi$, which is called the regularizer. We allow the regularizer $\psi_{t}$ to be time-varying in this paper.
@@ 5       The receiver calculates ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$;
The protocol of the proposed algorithm is illustrated in Algorithm 1. In the following, we establish the theoretical guarantee of the proposed algorithm via offering an upper bound on the expected error probability with learned channel decoders $\{\mathbf{G}_{t}\}_{t\in[T]}$.
@@
Theorem 1 provides a performance guarantee of the learned channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$: if $\exists\mathbf{G}^{*}\in\mathcal{G}$ such that the term $\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ is sub-linear w.r.t $T$, i.e., $\sum_{t\in[T]}\tilde{\ell}_{t}(G^{*})=\mathcal{O}(T^{\alpha}),\alpha<1$, the average expected error probability satisfies $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(\mathbf{G}_{t})\rightarrow 0$ as $T\rightarrow\infty$. This is because $\mathbb{E}\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}\leq 4L^{2}$, and we have $\frac{1}{T}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}}\leq\frac{2ML\sqrt{T}}{T}=\mathcal{O}(\frac{\sqrt{T}}{T})$. Besides, we notice that a larger positive auto-correlation function $\mathbb{E}[\langle\mathbf{H}_{t},\mathbf{H}_{t-1}\rangle]$ of $\{\mathbf{H}\}_{t\in[T]}$ results in a tighter derived upper bound. This indicates that the proposed method utilizes the channel correlation (second-order channel statistics) to improve the performance of learned decoder.
@@
Corollary 1 implies that the variance $\sigma^{2}_{W}$ of the channel noise $W$, the cumulative variance $(\sigma^{\mathbf{H}}_{1:T})^{2}$ and the cumulative variation $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ of channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ can deteriorate the performance of the learned channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$. We observe that smaller $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ leads to a tighter upper bound of average expected error probability $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(\mathbf{G}_{t})$, reflecting that the proposed method utilizes the distribution dependence of the time-correlated channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ to enhance the online optimization process.
@@ 1 Initialize: $\mathbf{w}^{\prime}_{1}=\arg\min_{\mathbf{w}\in\Omega}\psi_{1}(\mathbf{w})$ ;
$\quad+\frac{1}{T}\sqrt{\sum_{t\in[T]}\sum_{i\in[N]}\frac{1}{M}\sum_{j\in[M]}\sigma[\mathbbm{1}_{t}^{j}(i)]}\Big{)},$

### v5

@@ Abstract
Machine learning techniques have garnered great interest in designing communication systems owing to their capacity in tackling with channel uncertainty. To provide theoretical guarantees for learning-based communication systems, some recent works analyze generalization bounds for devised methods based on the assumption of Independently and Identically Distributed (I.I.D.) channels, a condition rarely met in practical scenarios. In this paper, we drop the I.I.D. channel assumption and study an online optimization problem of learning to communicate over time-correlated channels. To address this issue, we further focus on two specific tasks: optimizing channel decoders for time-correlated fading channels and selecting optimal codebooks for time-correlated additive noise channels. For utilizing temporal dependence of considered channels to better learn communication systems, we develop two online optimization algorithms based on the optimistic online mirror descent framework. Furthermore, we provide theoretical guarantees for proposed algorithms via deriving sub-linear regret bound on the expected error probability of learned systems. Extensive simulation experiments have been conducted to validate that our presented approaches can leverage the channel correlation to achieve a lower average symbol error rate compared to baseline methods, consistent with our theoretical findings.
@@ I Introduction
Most existing studies utilizing machine learning approaches to design communication systems lack theoretical justification for their proposed methods, and typically regard learned communication systems as black boxes [5, 6]. Recently, a few works are aiming to conduct theoretical analyses for learning-based communication systems [7, 8, 9]. Specifically, these studies leverage statistical learning theory [10, 11] to derive generalization bounds based on the Independently and Identically Distributed (I.I.D.) channel assumption. Building upon this assumption, the generalization capacity of learning-based communication systems can be assured. This is because these systems can be trained using sufficient data during the offline stage and demonstrate commendable performance when evaluated with data sampled from the same distribution.
However, meeting the assumption of I.I.D. channels in reality is generally challenging, given that practical communication scenarios often involve time-varying channels with dynamic statistical properties. Among time-varying channels, a prominent example is the time-correlated channel [12]. In real-world communication processes, there exist numerous instances of time-correlated channels. For example, user mobility usually leads to time-correlated fading channels in mobile communication [13, 14]. Specifically, in vehicular ad-hoc network scenarios (VANET), wireless links between moving vehicles and roadside units (RSUs) are significantly affected by vehicles’ mobility [15]. As a consequence of moderate user mobility, channel fading gains become interdependent between consecutive time slots, resulting in the well-known Markov fading channels [16, 17]. Notably, the nonlinear behavior of hardware components transforms additive white noise in the receiver into colored noise, thereby introducing temporal correlation [18]. Besides, time-correlated channel noise can arise due to narrowband filtering of uncorrelated impulsive noise by the RF front-end [19]. In these scenarios, establishing generalization bounds for learning-based communication systems over time-correlated channels is challenging. This is because classic tools like concentration inequalities from statistical learning theory [20], which heavily depend on the strong I.I.D. channel assumption, are not applicable in these scenarios.
@@
In this paper, we will establish online optimization algorithms [21, 22] for learning communication systems over time-correlated channels with solid theoretical guarantees. Specifically, we consider optimizing the uplink communication process between a mobile user equipment acting as the transmitter and a base station serving as a receiver. One potential application of proposed algorithms is to enhance the performance of distributed learning conducted by multiple mobile devices within wireless networks [23, 24]. Notably, our framework directly supports learning the downlink component of communication systems by enabling users to execute online optimization algorithms locally. In each communication round, we consider that the transmitter sends a package to the receiver. This package is composed of two parts: the first part is a deterministic training sequence known to both the transmitter and receiver, and the second part consists of data symbols that carry the actual information intended for transmission. These data symbols are random and unknown to the receiver. We highlight that this training sequence serves a similar role of pilot signals used in channel estimation in mobile communication, which is essential for calculating the loss required for conducting online optimization. The outlined online optimization procedure is depicted in Fig. 1. Following the assumption widely adopted in the study on using online learning to optimize communication systems [25, 26], we presume that the channel remains constant during the transmission of both the training sequence and the data symbols within a single communication round. While the channel statistics varies across different rounds. Consequently, channel decoders and codebooks learned by training sequences can be utilized to improve the transmission performance of data symbols within the same round. The transmission process of uplink communication is shown in Fig. 2.
@@
For the time-correlated channels studied in this paper, to be specific, we focus on two types of channels: 1) time-correlated fading channel with additive white Gaussian noise, where the fading distribution is unknown and time-correlated; and 2) time-correlated additive noise channel with an unknown and time-correlated noise distribution. For the time-correlated fading channel, the transceiver is equipped with a fixed codebook (constellation) and is tasked with optimizing the channel decoder using collected channel input-output pairs data. The optimized decoder processes channel outputs to compensate for fading-induced distortions—similar to channel equalization—thus enhancing symbol detection and overall transmission performance. Besides, based on the proposed convex surrogate loss, we can regard such online channel decoder learning problem as an online convex optimization problem. In the context of time-correlated additive noise channels with deterministic identity channel gains, the primary task of transceivers shifts to optimizing codebook selection to mitigate the effects of channel noise and enhance overall transmission performance. Similar to [7], the codebook considered in this paper essentially refers to a modulation scheme, i.e., a constellation. We model this online codebook learning problem as a multi-armed bandit problem: the transceiver selects optimal codebooks from a predefined super-codebook using empirical data collected in real-time.
@@ II Related works
In this section, we begin by presenting recent studies that have utilized machine learning techniques in the design of communication systems. Next, we provide an overview of research in online optimization, with emphasis on the application of optimistic OMD on predictable benign environments.
As previously mentioned, the remarkable success of machine learning algorithms has spurred interest in their application to optimize communication systems [5]. Some efforts focus on replacing conventional components of communication systems with modules that are learned from empirical data [30]. Notable examples include [31] for channel estimation and equalization, and [32] for channel coding and decoding. Another approach enhances traditional algorithms by integrating deep neural networks [33]. For example, in [34], DNNs are applied to enhance the Viterbi algorithm, and in [35], they are used for improving the performance of belief propagation. On the other hand, some recent works leverage statistical learning theory for establishing generalization bounds of learned communication systems. For instance, [7] utilize channel input-output pair data to optimize channel decoders and constellations, and derive generalization bounds for learned communication schemes using Rademacher complexity. [8, 9] employ the probably approximately correct (PAC) learning framework to provide theoretical guarantees for learned communication systems under discrete memoryless channels. Besides, the authors in [36] derive generalization bounds for their proposed neural belief propagation decoders. [37] leverage the margin theory to provide theoretical guarantees for learning-based deocders. Nonetheless, previous works have predominantly conducted a theoretical analysis based on the I.I.D. channel assumption, which is not applicable to with most practical communication scenarios.
Online optimization can effectively model numerous online machine learning problems, encompassing online convex optimization [21], prediction with expert advice [38], multi-armed bandit [39, 40] and more. This framework deals with a sequential decision problem, where a learner repeatedly takes actions within a feasible set and encounters a potentially adversarial loss function from the environment [21, 22]. During this decision-making process, the learner endeavors to devise an algorithm that minimizes regret, defined as the disparity between the total loss incurred by the learner and that of the best decision in hindsight. Departing from statistical learning theory, online optimization theory can furnish theoretical guarantees for algorithms without relying on the I.I.D. assumption. Recently, it has been noted that practical environments often exhibit predictable patterns, which the learner can leverage to achieve reduced regret [27, 29]. To pursue this objective, an optimistic Online Mirror Descent (OMD) framework has been proposed to harness environmental correlations for improving the performance of learned decisions [27, 29]. Furthermore, this framework is utilized to solve the semi-bandit problem in predictable environments [28] and the authors in [41] leverage it to address the stochastically extended adversarial online convex optimization problem.
@@ III-B Time-varying Channel Models
In this paper, we focus on two specific cases of this time-varying channel model, both of which hold high relevance in realistic scenarios. The first case involves a time-correlated fading channel, assuming that $f_{t}$ is a time-correlated linear transformation while the channel noise $Z_{t}$ is an I.I.D. Gaussian noise. The second case pertains to a time-correlated additive noise channel, where $f_{t}$ is an identity transformation and $Z_{t}$ is a time-correlated additive channel noise.
@@ IV-A Preliminaries
where $r\geq 1$. (a) holds based on the Boole’s inequality. (b) follows from the fact that $\mathbb{E}[\mathbbm{1}\{\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}-\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\leq 0\}]\leq\mathbb{E}\big{[}r-\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}\big{]}_{+}$ when $r$ satisfies $r\geq 1$, which can be found in [38].
@@ IV-B Optimistic OMD with Euclidean Regularizer
where $\mathcal{B}_{\psi}(X,Y)=\psi(X)-\psi(Y)-\langle\nabla\psi(Y),X-Y\rangle$ denotes the Bregman divergence induced by a differentiable convex function $\psi$, which is called the regularizer. We allow the regularizer $\psi_{t}$ to be time-varying in this paper. In essence, by performing twice mirror descent, our algorithm exploits channel correlation to enhance online optimization processes.
@@ 5       The receiver calculates ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$;
The protocol of the proposed algorithm is illustrated in Algorithm 1. In practice, we can adaptively modify the filter coefficients of a linear discrete-time filter according to the $\mathbf{G}_{t}$ learned by the optimistic OMD algorithm. Since adaptive equalization is a widely used technique in modern communication systems, our algorithm can be feasibly deployed in real-world systems. In the following, we establish the theoretical guarantee of the proposed algorithm via offering an upper bound on the expected error probability with learned channel decoders $\{\mathbf{G}_{t}\}_{t\in[T]}$.
@@
Theorem 1 provides a performance guarantee of the learned channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$: if $\exists\mathbf{G}^{*}\in\mathcal{G}$ such that the term $\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ is sub-linear w.r.t $T$, i.e., $\sum_{t\in[T]}\tilde{\ell}_{t}(G^{*})=\mathcal{O}(T^{\alpha}),\alpha<1$, the average expected error probability satisfies $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(\mathbf{G}_{t})\rightarrow 0$ as $T\rightarrow\infty$. This is because $\mathbb{E}\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}\leq 4L^{2}$, and we have $\frac{1}{T}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}}\leq\frac{2ML\sqrt{T}}{T}=\mathcal{O}(\frac{\sqrt{T}}{T})$. Moreover, we emphasize that this upper bound is meaningful when $\mathbb{P}_{t}(\mathbf{G}_{t})\leq\mathbb{E}\tilde{\ell}_{t}(\mathbf{G}^{*}),\forall t\in[T]$, as in this case $\frac{1}{T}\sum_{t}\mathbb{P}_{t}(\mathbf{G}_{t})\leq\frac{1}{T}\sum_{t}\mathbb{E}\tilde{\ell}_{t}(\mathbf{G}^{*})$, which vanishes as $T\rightarrow\infty$.
We notice that a larger positive auto-correlation function $\mathbb{E}[\langle\mathbf{H}_{t},\mathbf{H}_{t-1}\rangle]$ of $\{\mathbf{H}\}_{t\in[T]}$ results in a tighter derived upper bound. This indicates that the proposed method utilizes the channel correlation (second-order channel statistics) to improve the performance of learned decoder. For the practical implementation of our method, we do not need to tune $r_{t}$ based on the prior knowledge of $L$ and $D$. Instead, we compute the sub-gradient as ~$\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})=\frac{2}{M}\sum_{j\in[M]}\sum_{j^{\prime}\neq j}(\mathbf{x}^{j^{\prime}}-\mathbf{x}^{j})(\mathbf{y}_{t}^{j})^{T}$ used to update decoders, corresponding to set $r_{t}=2d^{*}DL+\frac{1}{\sqrt{T}}$.
@@
Corollary 1 implies that the variance $\sigma^{2}_{W}$ of the channel noise $W$, the cumulative variance $(\sigma^{\mathbf{H}}_{1:T})^{2}$ and the cumulative variation $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ of channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ can deteriorate the performance of the learned channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$. We observe that smaller $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ leads to a tighter upper bound of average expected error probability $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(\mathbf{G}_{t})$, reflecting that the proposed method utilizes the distribution dependence of the time-correlated channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ to enhance the online optimization process. Based on Assumption 1, we know that this upper bound becomes $\mathcal{O}\left(\frac{\gamma_{H}M\sqrt{T}}{T}+\frac{M\sqrt{T}}{T}\sqrt{\sigma_{W}^{2}}\right)$, which vanishes as $T\to\infty$.
Lastly, we briefly discuss practical applications of our method. For example, in MIMO systems with high-mobility users, our proposed optimistic OMD method enables dynamic precoding matrix updates via pilot signals, significantly enhancing overall transmission performance.
@@ 1 Initialize: $\mathbf{w}^{\prime}_{1}=\arg\min_{\mathbf{w}\in\Omega}\psi_{1}(\mathbf{w})$ ;
$\quad+\frac{1}{T}\sqrt{\sum_{t\in[T]}\sum_{i\in[N]}\frac{1}{M}\sum_{j\in[M]}\sigma[\mathbbm{1}_{t}^{j}(i)]}\Big{)}$
$\leq\tilde{\mathcal{O}}(\sqrt{\frac{N}{T}}),$